The bit was run to 4628 m, and the APC coring assembly, fitted with an APCT temperature shoe, was deployed. After a bottom-water temperature measurement had been taken a few meters above the seafloor, Hole 1089A was spudded at 0450 hr on 20 December with a core that recovered 7.3 m of sediment and fixed the seafloor depth at 4630.2 mbrf. Core orientation began with Core 4H.
Operations at Hole 1089A were plagued by numerous problems. Cores 3H, 4H, and 6H suffered core liner failures, with reduced recovery and disturbed core. The attempted APCT measurement with Core 4H was unsuccessful due to heave motion. Another downhole temperature was attempted with Core 7H. After a 10-min equilibration period, the overshot shearpin beneath the Tensor pressure case had failed, so a second run was required to retrieve the APC assembly. The APC was engaged successfully, but the coring line parted above the sinker bars. The remaining wire was reeled onto the winch, and the broken end was found about 125 m short of the ropesocket. A wireline fishing operation retrieved the broken wire and sinker bar assembly, and a pipe trip was avoided. It was discovered that the overshot shearpin had failed when the line parted and the APC was dropped, so it then became necessary to make a third wireline run to retrieve the APC.
Because of the two coring line failures in quick succession, the new forward wire was put into service. The aft wire was relegated to backup status.
When Core 7H arrived on deck, it was found to have a completely collapsed liner, and the APCT record was again degraded by vessel motion. APCT runs were discontinued at that point. Coring proceeded with liner failures of varying severity on nearly every core. About every other core had dismal recovery. Vessel heave, as noted by the driller, reached 12 ft at times. Coring conditions improved on the morning of 21 December. Core liner failures persisted, but generally were not catastrophic and had less of a disruptive effect on the stiffer sediment.
Hole 1089A was terminated at a total depth (TD) of 216.3 mbsf, short of APC refusal, after Core 23H. The decision to abandon was primarily made to take advantage of good weather conditions for coring the upper sediment section in Hole 1089B. The drill string was pulled above the seafloor at 1315 hr on 21 December to begin operations at Hole 1089B.
A stratigraphic overlap was desired, so the initial APC core interval was positioned 3.5 m higher than the equivalent interval of Core 1089A-1H. The spud attempt yielded no sediment and a broken core liner. Two additional "mudline" cores were attempted from 1 m deeper than the first attempt. In both cases the core liners broke near the top and the core catchers contained plastic liner fragments and only traces of sediment. On the fourth spud attempt the bit was lowered to 4630 mbrf in an effort to recover more consolidated sediment that would not be so easily washed from the core barrel while still maintaining an overlapping depth section. That attempt yielded a core with 4.8 m of sediment (and a broken liner). On the basis of recovery, seafloor depth was set at 4634.7 mbrf, but there is a strong possibility that some core was lost, which would make the seafloor depth too deep.
Coring success improved greatly on 22 December. Every effort was made to minimize core liner failures by adopting measures to reduce shock and stress on the brittle liners. Those included elimination of APCT runs and core orientation, trimming down APC piston seals, closing more APC speed ports, etc. It is not clear, though, whether these measures were responsible for the dramatic improvements, given the circumstances of similar coring problems later in the cruise. Some liner failures persisted, but they had a minor effect on core recovery and quality. Continuous APC cores were taken to a depth of 265 mbsf. The final two cores, 28H and 29H, exhibited incomplete stroke and recovery and about 40 kips overpull. A final core attempt resulted in a misrun when the APC apparently didn't land and seal properly at the outer core barrel. Hole 1089B then was terminated because the scientific objectives had been achieved.
Excellent weather and minimal vessel heave conditions continued as a partial seafloor core was attempted with the core bit at 4625 mbrf. The first core recovered 2.4 m of sediment, setting the seafloor depth at 4632.1 mbrf. Core 2H was started 2 m below the bottom of Core 1H to provide stratigraphic overlap. Continuous coring then proceeded, with orientation beginning on Core 4H. Good recovery was achieved, but core liner failures were experienced on about 50% of the cores. Most of the failures were small implosions near the top of the liner that had little effect on the core.
Weather conditions deteriorated through the day with wind gusts up to 32 kt, rain showers, and a near reversal of wind direction in the afternoon. Two sets of 3-m swells arrived at nearly right angles. As a result, the vessel could not take a heading to minimize roll and heave. Rig operating limits were approached, with rolls of 9° and heaves of up to 14 ft. During the period of increasing vessel motion, double wireline trips were required for Cores 11H (did not actuate) and 13H (sheared overshot pin and mechanical actuation). Conditions improved dramatically in the late evening as the wind shifted and dropped to a light breeze. One set of swells died rapidly, and operating conditions improved just as quickly. Coring was halted at 194.4 mbsf when scientific objectives were reached.
A mudline core fixed the seafloor depth at 4628.5 mbrf and continuous APC cores were taken to a TD at 118 mbsf. Again orientation began with Core 4H. Excellent core recovery and sediment conditions were achieved under good operating conditions. Only two minor core liner failures occurred in the 13 cores recovered.
The fourth penetration at Site 1089 completed stratigraphic coverage, filling all recovery gaps in the upper portion of the section. With the scientific coring objectives for the site attained, coring ceased at 1900 hr on Christmas Eve. The drill string was tripped, and the vessel got under way at 0530 hr on 25 December.
To 177 Operations - Site 1090
To 177 Table of Contents